Calculate the amount of heat liberated (in kJ) from 394 g of mercury when it cools from 57.2 C to 12.0 C. The specific heat of mercury is 0.139 J/g C.
1
Expert's answer
2015-07-03T03:32:47-0400
The heat absorbed or released is equal to the mass of the subject · its specific heat · the change in temperature:
q = m · s · Δt
So, the heat liberated is:
q = 394 · 0.139 ·(57.2-12.0)=2475.4 J, or 2.4754 kJ

Answer: 2.4754 kJ